OpenStudy (anonymous):

Can someone explain to me how to solve this problem and answer.... Martin’s average score after 4 tests is 89. What score on the 5th test would bring Martin’s average up to exactly 90?

OpenStudy (anonymous):

94

OpenStudy (amistre64):

lets see what the problem states; and see if we can make sense of it. we should be aware of how to find an average; its the sum of the parts divided by how many there are.... right?

OpenStudy (anonymous):

(4*89+x)/5 = 90 x=94

OpenStudy (amistre64):

\[\frac{sum.of.4.scores}{4}=89\] \[sum.of.4.scores=89(4)=356\] \[\frac{356+new.score}{5}=90\] \[356+new.score=90(5)\] \[new.score=90(5)-356\]

OpenStudy (anonymous):

thank you prashant :)